Calculate the mean,median, and the mode for the following data set:
Also find the standard deviation

5,5,5,5,5,10
1
Expert's answer
2013-02-18T10:04:55-0500
The mean value:
X = (5+5+5+5+5+10)/6 = 35/6 = 5.83

The median is the middle value in a set of values.:
5,5,5,5,5,10
Therefore median = 5.

The mode of a set is the value or values that occur most frequently. The number 5 occurs 5 times, and no other numbers occur as frequently. Therefore, the mode = 5

The standard deviation:
sqrt[ (5*(5-5.83)^2 +(10-5.83)^2)/(6-1) ] = 2.04